OpenStudy (anonymous):

3/n-1+2/3=5/3 ?please help

OpenStudy (anonymous):

\[\frac{3}{n-1}+\frac{2}{3}=\frac{5}{3}\]??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

subtract \(\frac{2}{3}\) from both sides, get \[\frac{3}{n-1}=1\]

OpenStudy (anonymous):

and so if \(\frac{3}{n-1}=1\) we know \(n-1=3\) making \(n=4\)

OpenStudy (anonymous):

aww thank you :)

OpenStudy (anonymous):

yw